OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of one half.

OpenStudy (anonymous):

y − 6 = one half(x + 3) y = one halfx − 6 y + 3 = one half(x − 6) x − 2y = 12

OpenStudy (anonymous):

its b because thats the only one in point slope form lol.

OpenStudy (anonymous):

xD want to help with a few more?

OpenStudy (anonymous):

oh shoot, its actually c

OpenStudy (anonymous):

y – y1 = m(x – x1)

OpenStudy (anonymous):

I'm dumb. was thinking of slope-intercept form.

OpenStudy (anonymous):

y + 3 = one half(x − 6) is my answer then?

OpenStudy (anonymous):

exactly I'm 100% sure this time. lol

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

Given the equation y − 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in standard form.

OpenStudy (anonymous):

−three fourthsx + y = 10 3x − 4y = −40 y = three fourthsx + 12 y = three fourthsx + 10

OpenStudy (anonymous):

y-4=3/4(x+8) y=3/4(x+8)+4 4y-3x=36?

OpenStudy (anonymous):

say waaa?

OpenStudy (anonymous):

well anyway, would have to be b because is in standard form. Ax+By=C

OpenStudy (anonymous):

xD I am trying to solve it right now.

OpenStudy (anonymous):

hold on.

OpenStudy (anonymous):

Yeah, that's what I got, sort of. xD

OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −one fifth(x − 2) y − 2 = −one fifth(x + 3)

OpenStudy (anonymous):

y2-y1/x2-x1 1-2/2-(-3) -1/5 Is slope y=-1/5x+b 1=-1/5(2)+b 1=-.4b -2.5=b is y-int y=-1/5x-2.5

OpenStudy (anonymous):

I seem a bit dilapidated. xD

OpenStudy (anonymous):

They put it in standard form so 5y=-x-2.5 5y+x=2.5

OpenStudy (anonymous):

-y=5x-2.5 -y=2.5(2x-1) Igive up

OpenStudy (anonymous):

xD Thanks for your help.

OpenStudy (anonymous):

The equation is of the form y - y1 = m(x - x1) (x1, y1) is either of the given points; "m" is the slope. Find the slope. m = (y2-y1)/(x2-x1) ....= (1-2)/(2+3) = -1/5 However, this slope is not in any of the options, and I checked my work, so either you mistyped something, maybe one of the points, or the people making the book made a mistake. Plug in either point for y1 and x1 in the equation. Apparently they used (-3, 2). y - 2 = -(1/5)(x - (-3)) y - 2 = -(1/5)(x + 3)

OpenStudy (anonymous):

lol no one else can figure it out